OpenStudy (anonymous):

find the slope of 2x-3y=12

OpenStudy (anonymous):

Solve the equation for y 2x-3y=12 subtract 2x from both sides -3y=-2x+12 divide both sides by -3 y=2/3x-4 slope is 2/3

OpenStudy (the_fizicx99):

2x-3y=12 -2x -2x -3y = -2x + 12 /-3 /-3 /-3 y = 2/3x - 4 Slope is 2/3 Y intercept is -4